# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

### v4

@@ 4 Experiments
gpt-3.5-instruct ? 67.3 (63.1–71.5) 25 87.8 (85.9–89.6) 11
gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 27
gpt-4o-mini ? 72.4 (68.1–76.3) 21 79.8 (77.3–82.2) 23
gpt-4o ? 74.4 (69.9–78.6) 17 83.7 (80.1–86.9) 21
claude-3.5 ? 84.0 (79.3–88.4) 7 87.1 (83.9–90.2) 15
o1-preview ? 76.5 (69.4–84.3) 24
Table 1: Results for our experiments. We present model metadata alongside model results on both the Transducer and Sequence completion tasks. Each cell contains the mean performance across DFAs for the best-performing prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses. “N/A” is used whenever the model returned an invalid result at least 25% of the time. (IT = Instruction-Tuned, TR/SR = Transducer/Sequence Completion rank, the ordinal rank of the given model on the given task.)
@@ 5 Results
Main results for all tasks are presented in Table 1. For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29. We then report the mean across DFAs and 95% bootstrap confidence intervals.
@@ 5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find that $n$-GramS heuristics tend to outperform LLMs.
As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results. Specifically, we find that BasicS is the best prompt for gpt-4o-mini, but not gpt-4o, where the best performing prompt is Red-GreenS. We find that claude-3.5 is entirely unable to follow the sequence completion prompts BasicS and More-ExplS, and performs best at the COTS prompt.
Additionally, we find that in this task, code-specific open-source models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example. Overall, the relative performances of LLMs and prompts comport somewhat well to heuristics on which models and prompting strategies should work best. Nonetheless, LLMs underperform simple $n$-Gram heuristics.
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix E for details.
Model Basic More-Expl COT Red-Green
Sequence Completion
gpt-4o-mini 72.4 (68.1–76.3) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
gpt-4o 72.1 (65.9–78.2) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
claude-3.5 N/A N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
Transducer
gpt-4o-mini 79.8 (77.3–82.2) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
gpt-4o 83.7 (80.1–86.9) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
claude-3.5 86.9 (83.3–90.0) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
@@ 5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ baseline only achieving 96.4% accuracy. Comparisons are still be valid as all models see the same fraction of unsolvable instances.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o. Even within a model class we find no clear pattern: GPT-4o and o1-preview [^4] The particularly poor performance of o1-preview may be due to the model not supporting temperature 0. In other experiments, we found that GPT models with temperature 1 tended to perform poorly. See Appendix B for more details. are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. To investigate that this is not specific to the Basic prompt, we investigate other prompts for chat models. As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
We conclude that LLMs are unable to perform well on the DFA transducer inference task. This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
@@ 5.3 Comparison of Benchmarks
Figure 3: Transducer and sequence completion results plotted against each other. Points represent the mean over several DFAs and intervals represent 95% confidence intervals. Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline and proprietary models.
Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks. While at a high level, there is a positive correlation between the two, there are a few notable differences. For one, the Code models perform notably better than other open source models on Sequence Completion, but not on Transducer. Additionally, on Transducer, a ceiling on performance is observed, where LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
5.4 Case Study: Sum Modulo 3 DFA
Correct Incorrect
Total 100% 100%
a is no-op 70% 73%
1b and 1c lead to 0 47% 57%
2-periodic 30% 47%
3-periodic 13% 13%
2 red rooms 7% 10%
We investigate the transducer task on the DFA depicted in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. For this case study, we focus on the model/prompt combinations MB and CR, defined as
  • MB: mistral-nemo-minitron-8B/BasicT. Selected as it is the best performing combination overall.
  • CR: claude-3.5/Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 4b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix C for details.
[…]
@@ 6 Conclusion
Our findings highlight significant weaknesses in large language models’ ability to generalize to entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs. These results, combined with that of previous work demonstrating that large language models can quite accurately perform a variety of language tasks, suggests that LLMs solve language problems via a mechanism distinct from general language reasoning ability. Our use of n-gram baselines and next-token prediction tasks allows us to exclude the possibility that the issue is primarily related to LLMs’ lack of world modeling or any inherent limitations of next-token prediction models. We believe our results suggest that LLMs have learned individual models of particular languages, but not a general theory of language.
Interestingly, in our transducer experiments, LLMs consistently perform better by directly predicting the next token than by explicitly reasoning through the problem. While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
Many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages. More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format. For these tasks, it may be prudent to seek a new approach.
@@ A.3 Sampling of Transducer Tasks
Appendix B Results of o1-preview
We evaluated o1-preview on 10 DFAs, using 30 problem instances per DFA of the Transducer task, as in other Transducer experiments, and the $\textsc{Basic}_{T}$ prompt, as this is the most neutral prompt. Table 3 displays results on each DFA. Overall, while these results are not on a particularly large sample, they fairly definitively demonstrate that o1-preview does not achieve strikingly good performance on this task.
DFA o1-preview gpt-4o 6-Gram
1 25/30 27/30 26/30
2 23/29 24/30 25/30
3 19/30 23/30 28/30
4 22/30 23/30 28/30
5 29/29 30/30 30/30
6 19/30 24/30 30/30
7 17/29 23/30 25/30
8 23/30 25/30 26/30
9 21/30 28/30 30/30
10 29/30 29/30 30/30
Table 3: Results on each DFA. We find that in 7 cases, o1-preview underperforms gpt-4o, in 2 cases it gets the same number of instances wrong but provides a non-answer on an additional instance, and in 1 case it ties gpt-4o. In no cases does it outperform.
Appendix C Transducer results by difficulty class
@@
Appendix D More details on Sum Modulo 3 DFA case study
Figure 6 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions. In the (a) condition, the trace ends in such a way that a 3-Gram model would be able to determine the output, and the (b) condition is the complement.
@@ Appendix D More details on Sum Modulo 3 DFA case study
Appendix E Sequence Completion task prompt with Commas
@@
Appendix F Prompt Listings

### v6

@@ 4 Experiments
gpt-3.5-instruct ? 67.3 (63.1–71.5) 31 87.8 (85.9–89.6) 14
gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 32
gpt-4o-mini ? 72.4 (68.1–76.3) 26 79.8 (77.3–82.2) 28
gpt-4o ? 74.8 (69.3–80.4) 21 83.7 (80.1–86.9) 25
claude-3.5 ? 82.8 (77.5–87.5) 8 86.9 (83.3–90.0) 20
o3-mini ? 81.1 (76.0–85.8) 9 74.7 (70.7–78.8) 29
gpt-5 ? 77.9 (71.6–84.0) 15 83.6 (79.9–87.1) 26
Table 1: Results for our experiments on both the Transducer and Sequence completion tasks. Each cell contains the mean performance across DFAs for the best-performing Basic prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses. “N/A” is used whenever the model returned an invalid result at least 25% of the time. (IT = Instruction-Tuned, TR/SR = rank of the given model on Transducer/Sequence Completion.)
Model Basic Basic-COT More-Expl DFA-COT Red-Green
Sequence Completion
gpt-4o-mini 72.4 (68.1–76.3) 60.0 (55.8–64.4) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
gpt-4o 72.1 (65.9–78.2) 74.8 (69.3–80.4) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
claude-3.5 N/A 82.8 (77.5–87.5) N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
o3-mini N/A 81.1 (76.0–85.8) N/A 58.2 (49.6–66.8) 69.8 (64.4–75.0)
gpt-5 77.9 (71.6–84.0) 75.1 (68.6–81.7) 68.9 (61.0–76.7) 66.0 (58.9–72.8) 87.5 (83.1–91.5)
Transducer
gpt-4o-mini 79.8 (77.3–82.2) 66.5 (64.3–68.7) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
gpt-4o 83.7 (80.1–86.9) 67.8 (62.4–73.2) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
claude-3.5 86.9 (83.3–90.0) 74.2 (70.0–78.3) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
o3-mini 72.8 (68.4–77.3) 74.7 (70.7–78.8) 74.7 (70.3–79.2) 86.1 (83.9–88.4) 92.4 (91.3–93.5)
gpt-5 83.6 (79.1–87.7) 83.6 (79.9–87.1) 85.2 (81.0–88.8) 96.7 (95.3–98.0) 96.6 (95.4–97.8)
@@ 5 Results
Main results for all tasks are presented in Table 1. We ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29. We then aggregate across DFAs. All comparisons involving 4-Gram, 5-Gram, and 6-Gram to all other models are statistically significant (see Appendix H for details).
@@ 5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100% accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find that $n$-GramS heuristics tend to outperform LLMs.
As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results. Specifically, we find that for gpt-4o-mini, immediately predicting a next token seems to be better, while for gpt-4o and gpt-5 there is no large effect. claude-3.5 and o3-mini are unable to answer the BasicS prompt at all, but outperform other proprietary models when using Basic-COTS. In this task, revealing the problem structure appears to not have a massive effect on performance, with gpt-5 being the only model to incorporate this information into a statistically significantly improved performance, and even then only in the Red-GreenS word problem prompt (still underperforming 4-GramS).
Additionally, we find that in this task, code-specific open-weight models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example. Overall, the relative performances of LLMs and prompts generally comport to heuristics on which models and prompting strategies should work best (with the notable exception of gpt-5). Nonetheless, LLMs underperform simple $n$-Gram heuristics.
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. We found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix F.1 for details; we confirmed that the LLMs we investigate could reason about strings without commas, see Appendix F.2 for details.
@@ 5.2 Transducer
Unlike sequence completion, this task is not fully determined, with $\textsc{BruteForce}_{T}$ getting 96.4% accuracy. Comparisons are still valid as all models see the same fraction of unsolvable instances.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not correspond to their overall scale, with open weight LLama-3 and Mistral Nemo 8B parameter models outperforming much larger proprietary models. Even within a model class we find no clear pattern: all other GPT models are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. More specifically, as seen in Table 2, our Basic-COTT prompt results in underperformance by all non-reasoning models, suggesting that models are generally most able to solve this task when it is a simple next-token-prediction task. Providing the problem structure also does not help non-reasoning models improve substantially, but does allow o3-mini to perform well, and gpt-5 to completely solve the task (achieving parity with BruteForce), demonstrating that reasoning models’ underperformance at the original ICL language reasoning task is not due to the underlying difficulty of the task itself.
We conclude that LLM ICL is unable to perform well at language inference. This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar and unknown source, even a relatively simple one.
@@ 5.3 Comparison of Benchmarks
Figure 3: Transducer and sequence completion results plotted against each other. Points are the mean over several DFAs, with 95% confidence intervals. Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline & proprietary models.
Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks. While at a high level, there is a positive correlation between the two, there are a few notable differences. For one, the Code models perform notably better than other open weight models on Sequence Completion, but not on Transducer. Additionally, on Transducer, a ceiling on performance is observed, where non-reasoning LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
@@ 6 Conclusion
Our findings highlight significant weaknesses in large language models’ ability to in-context-learn entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs. These results, combined with that of previous work demonstrating that large language models can quite accurately perform a variety of language tasks, suggests that LLMs solve language problems via a mechanism distinct from general language reasoning ability. Our use of n-gram baselines and next-token prediction tasks allows us to exclude the possibility that the issue is primarily related to LLMs’ lack of world modeling or any inherent limitations of next-token prediction models. We believe our results suggest that LLMs have learned individual models of particular languages, but not a general theory of language.
Interestingly, in our transducer experiments, directly predicting the next token rather than explicitly reasoning through the problem works better, except for reasoning models where they perform similarly. While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
One potential goal for foundation models is to replace all machine learning with ICL. Our results suggest that current models are not progressing towards this goal.
@@ A.3 Sampling of Transducer Tasks
Appendix B Transducer results by difficulty class
@@
Appendix C Case Study: Sum Modulo 3 DFA
Correct Incorrect
Total 100% 100%
a is no-op 70% 73%
1b and 1c lead to 0 47% 57%
2-periodic 30% 47%
3-periodic 13% 13%
2 red rooms 7% 10%
We investigate the transducer task on the DFA depicted in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. For this case study, we focus on the model/prompt combinations MB (mistral-nemo-minitron-8B/BasicT: the best performing non-reasoning model combination overall) and CR (claude-3.5/Red-GreenT: the best performing non-reasoning model combination that provides an explanation, needed later for our qualitative analysis).
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 5b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any non-reasoning LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix B for details.
The model also attempts to identify periodic patterns, but identifies period-2 patterns more than period-3 patterns, despite knowing that there are three “rooms” (states). At no point in any of the 60 reasoning traces analyzed does it realize that this is a version of the Sum Modulo 3 DFA [^5] In fact, in none of the 1000 traces do the substrings “sum” or “mod” appear, except as a part of “assuming”, but it does show some glimmers of world modeling: in a few cases it correctly determines that there are two red rooms; however, this does not lead to further discoveries. It is not superior reasoning that leads to correct solutions, rather the correct examples are more likely to be ones that a 3-Gram model would infer correctly, i.e., those traces ending in a, 1b, or 1c, which occur cumulatively in $\frac{5}{9}$ of cases [^6] On the $\sim$$\frac{5}{9}$ of examples following this pattern, CR achieves 93.5%, to the 6-GramT’s 97.3%, and on the remaining $\sim$$\frac{4}{9}$, it achieves 43.8%, to the 6-GramT’s 60.7%. Detailed Venn diagrams on these conditions can be found in Figure 6..
@@ Appendix C Case Study: Sum Modulo 3 DFA
Appendix D Increased number of examples
Figure 7: Accuracy by number of examples, as we vary the number from 30 to 600 (note the log scale). $\infty$-Gram is a model that finds the longest sequence that matches the ending of the sequence and copies the following token.
We varied the number of examples parameter from 30 to 600, investigating specifically mistral-nemo-minitron-8B as it is a better performing model on the transducer task. We find that for the Sequence Completion task, the $n$-Gram models do not generally improve as the number of examples increases, while the LLM does; however, the LLM remains below 4-Gram performance at all points. In the Transducer experiment, both the n-Gram models and the LLM improve, with the LLM crossing 4-Gram and 5-Gram performance and eventually roughly matching 6-Gram performance. However, there is now an increased gap between 6-Gram and 7-Gram that did not exist at 30 (we did not include 7-Gram or above in the main table for this reason). Overall, it is possible that the larger number of instances of the “correct n-Gram” appearing (i.e., the suffix of the prompt followed by the correct answer) causes the model to be better at producing predictions. As the LLM at no point outperforms the best n-Gram, we do not believe that this alters our overall conclusions.
Appendix E Nonzero temperature
Model Prompt Zero Temp Nonzero Temp Difference
Sequence Completion
mistral-nemo-minitron-8B Basic 78.70% (77.49% – 79.79%) 77.67% (76.51% – 78.76%) -1.04% (-1.49% – -0.63%)
claude-3.5 COT 84.00% (79.33% – 88.44%) 84.22% (79.56% – 89.00%) 0.22% (-2.11% – 2.33%)
claude-3.5 Red-Green 80.00% (74.89% – 85.22%) 80.78% (75.00% – 86.11%) 0.78% (-2.11% – 3.44%)
Transducer
mistral-nemo-minitron-8B Basic 88.56% (88.05% – 89.08%) 88.17% (87.64% – 88.68%) -0.39% (-0.58% – -0.22%)
claude-3.5 Basic 86.89% (83.33% – 90.00%) 87.00% (83.33% – 90.11%) 0.11% (-0.67% – 0.89%)
claude-3.5 More-Expl 87.11% (83.88% – 90.22%) 86.89% (83.22% – 90.11%) -0.22% (-1.11% – 0.67%)
claude-3.5 COT 76.44% (72.89% – 79.89%) 78.11% (74.66% – 81.33%) 1.67% (-0.56% – 3.78%)
claude-3.5 Red-Green 82.89% (78.89% – 86.89%) 82.78% (78.89% – 86.33%) -0.11% (-2.00% – 1.78%)
Table 3: Results varying temperature. Second column is a temperature of 0.1, third column is differences. In all columns, we annotate a 95% confidence interval, using paired differences for the third column.
In order to determine whether a small nonzero temperature might lead to better results, we investigated using a temperature of 0.1 for mistral-nemo-minitron-8B and claude-3.5 (these two chosen for the reasons described in Section C, notably o3-mini/gpt-5 already are using nonzero temperatures). We find that a temperature of 0.1 does not significantly change performance, resulting in no significant change for any model/task/prompt combination. The largest improvement we observe is claude-3.5 on COT on the Transducer task, which gains 1.7%, from 76.4% to 78.1% (still not enough to make it the best prompt for claude-3.5).
Appendix F Tokenization
@@
F.1 Sequence Completion task prompt with Commas
@@
F.2 Directly confirming models can read sequences of letters
To fully exclude the possibility that models are unable to read sequences of letters, we perform the following experiment: we take the regular expression ab(abc)+ and directly provide it to the model, then ask the model to test it on a string (one string provided per query), using the prompt I will give you a string. Tell me whether it matches the following regular expression: ’^ab(abc)+\$’ (without quotes). Just answer YES or NO. on one line, followed by a string on the next. We sample 100 random strings of length 2-17, via the following procedure (1) sample a random valid string uniformly (2) with 50% probability, randomly mutate one of the elements of the string to a different character. All strings are thus either correct or near-correct. We also add Answer (YES or NO): on a third line for non-chat models to encourage them to provide a response rather than another string.
Model Accuracy Non-response
llama3-8B 66%
llama3-70B 80% 26%
llama3.1-8B-Instruct 83%
llama3.1-8B 89%
llama3.1-70B 86% 36%
qwen-2.5-7B 62%
qwen-2.5-32B 100%
mistral-nemo-instruct-12B 99%
gemma-7b 67% 15%
starcoder2-15b 62%
codestral-22B 95%
qwen-2.5-coder-7B 100%
qwen-2.5-coder-instruct-7B 56%
qwen-2.5-coder-instruct-32B 96%
gpt-3.5-instruct 67%
gpt-3.5-chat 82%
gpt-4o-mini 100%
gpt-4o 100%
claude-3.5 100%
o3-mini 100%
gpt-5 100%
Table 5: Results on Regex task. As in the rest of this paper, accuracies are computed ignoring non-response. Models mistral-nemo-minitron-8B, mistral-nemo-base-12B, falcon-7b, deepseek-coder-33b-instruct have non-response rates over 98%.
Results for this are presented in Table 5. While not all models perform well at this task, performances follow roughly what one might expect from standard benchmarks (rather than the Sequence Completion results in the paper). Frontier proprietary models perform extremely well, larger open weight models tend to perform well, and smaller open weight models are more hit-and-miss (some performing well, some poorly).
The proprietary models newer than the GPT-3 series get 100% accuracy on this task, while performance is lower for the open weight models. There is no obvious relationship between models that perform well at this particular task and models that perform well at Sequence Completion, indicating that to whatever degree models are performing poorly at this test task, it is not because they are unable to process the string. Notably, o3-mini, gpt-4o, and gpt-4o-mini all perform fairly poorly at the Sequence Completion task, placing below median, but all achieve 100% on this task.
Open weight models are more mixed, with larger ones tending to do nearly as well as newer proprietary models (notable exception being llama3-70B), but smaller ones occasionally performing well and occasionally performing poorly.
Appendix G Model non-answers
Model Basic Basic-COT More-Expl DFA-COT Red-Green
Sequence Completion
llama3-8B 0.0 (0.0–0.0)
llama3-70B 0.0 (0.0–0.0)
llama3.1-8B-Instruct 0.0 (0.0–0.0)
llama3.1-8B 0.0 (0.0–0.0)
llama3.1-70B 0.0 (0.0–0.0)
mistral-nemo-minitron-8B 0.0 (0.0–0.0)
mistral-nemo-base-12B 0.0 (0.0–0.0)
mistral-nemo-instruct-12B 0.0 (0.0–0.0)
gemma-7b 0.0 (0.0–0.0)
falcon-7b 0.0 (0.0–0.0)
starcoder2-15b 0.0 (0.0–0.0)
codestral-22B 0.0 (0.0–0.0)
deepseek-coder-33b-instruct 0.0 (0.0–0.0)
qwen-2.5-coder-instruct-7B 0.0 (0.0–0.0)
qwen-2.5-7B 0.0 (0.0–0.0)
qwen-2.5-32B 0.0 (0.0–0.0)
qwen-2.5-coder-instruct-32B 0.0 (0.0–0.0)
qwen-2.5-coder-7B 0.0 (0.0–0.0)
gpt-3.5-instruct 2.5 (1.9–3.0)
gpt-3.5-chat 99.9 (99.7–100.0)
gpt-4o-mini 0.0 (0.0–0.0) 0.3 (0.2–0.5) 0.0 (0.0–0.0) 1.0 (0.6–1.4) 0.2 (0.1–0.4)
gpt-4o 4.4 (3.3–5.7) 1.9 (1.1–2.8) 100.0 (100.0–100.0) 5.0 (3.8–6.2) 8.4 (6.6–10.2)
claude-3.5 99.7 (99.2–100.0) 0.1 (0.0–0.3) 97.8 (96.9–98.6) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
o3-mini 80.2 (78.0–82.3) 3.8 (2.6–5.2) 91.6 (89.9–93.2) 5.7 (4.1–7.2) 0.4 (0.0–1.0)
gpt-5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 8.2 (4.7–12.1) 5.2 (3.1–7.4) 0.4 (0.1–0.9)
Transducer
llama3-8B 0.0 (0.0–0.0)
llama3-70B 0.0 (0.0–0.0)
llama3.1-8B-Instruct 0.0 (0.0–0.0)
llama3.1-70B 0.0 (0.0–0.0)
llama3.1-8B 0.0 (0.0–0.0)
starcoder2-15b 0.0 (0.0–0.0)
codestral-22B 0.0 (0.0–0.0)
deepseek-coder-33b-instruct 0.0 (0.0–0.0)
qwen-2.5-coder-7B 0.0 (0.0–0.0)
qwen-2.5-coder-instruct-7B 0.0 (0.0–0.0)
qwen-2.5-7B 0.0 (0.0–0.0)
qwen-2.5-32B 0.0 (0.0–0.0)
qwen-2.5-coder-instruct-32B 0.0 (0.0–0.0)
mistral-nemo-minitron-8B 0.0 (0.0–0.0)
mistral-nemo-base-12B 0.0 (0.0–0.0)
mistral-nemo-instruct-12B 0.0 (0.0–0.0)
gemma-7b 0.0 (0.0–0.0)
falcon-7b 0.0 (0.0–0.0)
gpt-3.5-instruct 0.0 (0.0–0.1)
gpt-3.5-chat 0.1 (0.0–0.3)
gpt-4o-mini 1.8 (1.3–2.3) 0.0 (0.0–0.0) 5.8 (4.8–6.9) 0.0 (0.0–0.0) 0.7 (0.4–1.0)
gpt-4o 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
claude-3.5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
o3-mini 0.1 (0.0–0.3) 0.0 (0.0–0.0) 0.4 (0.1–0.9) 0.0 (0.0–0.0) 0.0 (0.0–0.1)
gpt-5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.7 (0.1–1.4)
Table 6: Model non-answers, as a percentage of all prompt responses. A non-response is not included in accuracy computations for Table 1 or Table 2, but whenever it rises above 25%, N/A is placed in those tables.
Table 6 depicts the percentage of model non-answers by model and prompt. In general, this distribution is highly bimodal, with values always being either below 9% or above 97%.
The only prompt-vs-prompt orderings that are changed by scoring non-answers as 0 are that, on Sequence Completion, BasicS rises above Red-GreenS for gpt-4o, making it the best prompt; and that on Transducer, Red-GreenT for gpt-4o-mini rises above More-ExplT (though still behind BasicT. The qualitative conclusions about next token prediction vs chain of thought results remain the same.
The only change to relative model ordering is that on Sequence Completion, gpt-4o drops 8 ranks, from 17th place to 25nd place, being passed by several open weight models, gpt-4o-mini, and o3-mini. No change occurs on the transducer results. Qualiative conclusions about model ordering remain the same.
Appendix H Significance
Figure 8: Significance of comparisons between rows of Table 1. Results in blue and orange are significant, results in pink are not.
Figure 8 shows which comparisons between rows of Table 1 are significant. Significance computations are performed by running a 2-tailed bootstrap significance test on paired (by DFAs) differences.
Appendix I Prompt Listings